# Confidential Computing on nVIDIA H100 GPU: A Performance Benchmark Study

### v1

@@ Abstract
This report evaluates the performance impact of enabling Trusted Execution Environments (TEE) on NVIDIA H100 GPUs for large language model (LLM) inference tasks. We benchmark the overhead introduced by TEE mode across various models and token lengths, focusing on the bottleneck caused by CPU-GPU data transfers via PCIe. Our results show that while there is minimal computational overhead within the GPU, the overall performance penalty is primarily due to data transfer. For most typical LLM queries, the overhead remains below 5%, with larger models and longer sequences experiencing near-zero overhead.
@@ 1 Introduction
Trusted Execution Environments (TEEs), are increasingly important in machine learning and AI due to rising security requirements across both enterprise and decentralized applications. The introduction of TEE-enabled GPUs, such as the NVIDIA H100, adds an extra layer of protection for sensitive data but may affect performance. Understanding these trade-offs, especially for large-scale machine learning tasks, is crucial for adopting TEE in high-performance AI applications.
This report quantifies the performance overhead of enabling TEE on the NVIDIA H100 GPU during LLM inference tasks, examining where the overhead arises and under what conditions it can be minimized.
@@ 2.1 Trusted Execution Environment
A TEE is a hardware-based security feature that isolates computations, preventing unauthorized access and tampering—even from the operating system or the physical hardware owner. As the core technology to enable Confidential Computing, TEEs create secure enclaves where sensitive data and code are processed with encryption, ensuring confidentiality and integrity even if the broader system is compromised. Traditionally implemented in CPUs, TEE technology was extended to GPUs by NVIDIA in 2023, enabling tamper-proof and confidentiality-preserving computation inside the GPU with minimal performance penalty.
@@
The NVIDIA H100 GPU represents a significant milestone as the first GPU to support TEE. In TEE mode, the H100 operates in an isolated and secure environment where data transfers between the CPU and GPU are encrypted. This is facilitated by ”bounce buffers,” which protect all inputs and outputs during transit between the CPU’s encrypted memory and the GPU’s internal memory.
To maintain end-to-end security, the H100 works in conjunction with CPU TEEs, such as Intel’s TDX or AMD’s SEV-SNP, securing communication channels between the GPU driver and interacting software. This prevents unauthorized access and ensures data integrity throughout the process.
The H100 also implements remote attestation to verify the GPU’s identity and the authenticity of its firmware. Secure Boot further strengthens security by ensuring that only authenticated firmware is executed during the GPU’s boot process.
@@ 2.3 Performance Impact
Enabling TEE on the NVIDIA H100 GPU introduces performance overheads primarily due to additional encryption and decryption during secure data transfer. While the GPU’s internal computation remains unaffected, the main bottleneck lies in the CPU-GPU IO, particularly when data is exchanged via PCIe. This impact varies with the size of the data transfer. The following sections present experimental results quantifying these effects across various use cases.
@@ 3.1 Metrics
The primary metrics are evaluated following typical evaluation frameworks [AAK+24]:
@@
  • ITL (Inter Token Latency): The time between generating each token during decoding. This directly affects the perceived model speed. A rate of around 6 tokens per second is necessary for smooth user experience, assuming average reading speed.
  • TPS (Token per Second): The average rate of token generation during decoding. It’s calculated as the total decoding time divided by the number of tokens generated.
@@
  • QPS (Query per Second): The maximum load a system can handle while meeting latency targets. Higher QPS reduces serving costs and is a key measure of system capacity.
@@ 3.2 Test Scenarios
Experiments were structured to explore TEE mode impact under diverse conditions:
@@
  • Sequence Lengths: We tested various token lengths by sampling the ShareGPT Dataset [ano] to simulate different LLM inference tasks.
  • Batch Size: We tested both the fixed batch sizes (1, 4, and 16) and dynamic batch size determined by vLLM [KLZ+23] to simulate the performance for serving real-time requests and batch requests.
@@ 3.3.1 Infrastructure
We set up the experiments with the following hardware:
@@ 3.3.2 Application
The experiments utilized the Benchmark suites of vLLM v0.5.4 (rev: 4db5176) [KLZ+23].
@@ 4 Results
Conclusion 1: The average overhead is less than 7%. We quantified the overhead by measuring the throughput with TEE mode enabled versus disabled, across varying input sizes and model configurations as shown in Table 1.
@@
The throughput is measured in two ways: the average throughput of the outputted tokens per second (TPS), and that of the parallel requests the hardware can handle (PQS). TPS is measured by running the model with a batch size of 1. It shows the pure latency overhead introduced by the TEE mode. It reflects the performance of realtime requests. QPS is measured by maximizing the query throughput with dynamically optimized batch size. It reflects the minimal average overhead the TEE mode brings.
Conclusion 2: The overhead reduces toward zero as the model size grows. As shown in Table 1, the smallest model (LLama-3.1-8B) has the highest overhead. The medium size model (Phi3-14B-128k) has roughly 2/3 of the overhead compared with the smaller one. The largest model (Llama-3.1-70B) has a trivial overhead close to zero.
Conclusion 3: The latency is the main factor contributing to the overhead of the TEE model. Table 2 shows the overhead introduced to the latency measured by TTFT and ITL. TTFT has a higher overhead compared with ITL, indicating the bottleneck is likely introduced by the IO instead of the computation happening inside the TEE. Nevertheless, the overhead becomes trivial when hosting heavy computation models like Llama-3.1-70B.
@@
Conclusion 4: The overhead reduces as the token size grows. As shown in Figure 1, the throughput overhead reduces when the sequence length grows, measured by the total input and output token count. The detailed throughput metrics across various sequence length can be found at Table 3.
@@
Conclusion 5: TEE can reach a typical throughput
Our experiments revealed that, with medium-sized inputs, the NVIDIA H100 GPU achieves 130 TPS for Llama3-8B, while the larger Phi-3-14B model reaches approximately 6 TPS. These results demonstrate the robust performance of the H100 GPU across models of varying complexity.
@@ 5 Conclusion
Our results show that as input size grows, the efficiency of TEE mode increases significantly. When computation time within the GPU dominates overall processing time, the IO overhead introduced by TEE mode diminishes, allowing efficiency to approach nearly 99%.
Efficiency growth is more pronounced in larger models, such as Phi3-14B-128k and Llama-3.1-70B, due to their greater computational demands, which result in longer GPU processing times. Consequently, the IO overhead becomes increasingly trivial as model size increases.
The total token size (sum of input and output token size) significantly influences the throughput overhead. Larger total token counts lead to higher efficiencies, as they enhance the ratio of computation time to IO time.

### v2

@@ Abstract
This report evaluates the performance impact of enabling Trusted Execution Environments (TEE) on nVIDIA H100 GPUs for large language model (LLM) inference tasks. We benchmark the overhead introduced by TEE mode across various LLMs and token lengths, with a particular focus on the bottleneck caused by CPU-GPU data transfers via PCIe. Our results indicate that while there is minimal computational overhead within the GPU, the overall performance penalty is primarily attributable to data transfer. For the majority of typical LLM queries, the overhead remains below 5%, with larger models and longer sequences experiencing nearly zero overhead.
@@ 1 Introduction
Trusted Execution Environments (TEEs) are increasingly important in machine learning and AI due to growing security requirements in both enterprise and decentralized applications [SAB15, MSM+18, AKKH18]. The introduction of TEE-enabled GPUs, such as the nVIDIA H100, adds an extra layer of protection for sensitive data but may impact performance. Understanding these trade-offs, particularly for large-scale machine learning tasks, is crucial for adopting TEE in high-performance AI applications [YMY+22, WO24].
This report quantifies the performance overhead of enabling TEE on the nVIDIA H100 GPU during LLM inference tasks, identifying where the overhead arises and under what conditions it can be minimized.
@@ 2.1 Trusted Execution Environment
A TEE is a hardware-based security feature that isolates computations, preventing unauthorized access and tampering, even from the operating system or the physical hardware owner. As the core technology enabling Confidential Computing, TEEs create secure enclaves where sensitive data and code are processed with encryption, ensuring confidentiality and integrity even if the broader system is compromised [SAB15]. Traditionally implemented in CPUs, TEE technology was extended to GPUs by nVIDIA in 2023, enabling tamper-proof and confidentiality-preserving computation inside the GPU with minimal performance penalty [DGK+23].
@@
The nVIDIA H100 GPU marks a significant milestone as the first GPU to support TEE. In TEE mode, the H100 operates in an isolated and secure environment where data transfers between the CPU and GPU are encrypted. This is achieved through “bounce buffers”, which protect all inputs and outputs during transit between the CPU’s encrypted memory and the GPU’s internal memory [DGK+23].
To maintain end-to-end security, the H100 works in conjunction with CPU TEEs, such as Intel’s TDX [Int] or AMD’s SEV-SNP [AMD, SS20], securing communication channels between the GPU driver and interacting software. This setup prevents unauthorized access and ensures data integrity throughout the process.
The H100 also implements remote attestation to verify the GPU’s identity and the authenticity of its firmware. Additionally, Secure Boot ensures that only authenticated firmware is executed during the GPU’s boot process, further strengthening security.
@@ 2.3 Performance Impact
Enabling TEE on the nVIDIA H100 GPU introduces performance overheads primarily due to additional encryption and decryption during secure data transfer [MYF+24]. While the GPU’s internal computation remains unaffected, the main bottleneck lies in the CPU-GPU I/O, particularly when data is exchanged via PCIe. This impact varies with the size of the data transfer. The following sections present experimental results quantifying these effects across various use cases.
@@ 3.1 Metrics
The primary metrics were evaluated following typical evaluation frameworks [AAK+24]:
@@
  • ITL (Inter-Token Latency): The time between generating each token during decoding. This directly affects the perceived model speed. A rate of around 6 tokens per second is necessary for a smooth user experience, assuming an average reading speed.
  • TPS (Tokens per Second): The average rate of token generation during decoding. It is calculated as the number of tokens generated divided by the total decoding time.
@@
  • QPS (Queries per Second): The maximum load a system can handle while meeting latency targets. Higher QPS reduces serving costs and is a key measure of system capacity.
@@ 3.2 Test Scenarios
Experiments were structured to explore the impact of TEE mode under diverse conditions:
@@
  • Sequence Lengths: Various token lengths were tested by sampling the ShareGPT Dataset [ano] to simulate different LLM inference tasks.
  • Batch Size: Both fixed batch sizes (1, 4, and 16) and dynamic batch sizes determined by vLLM [KLZ+23] were tested to simulate the performance for serving real-time requests and batch requests.
@@ 3.3.1 Infrastructure
The experiments were set up with the following hardware:
@@ 3.3.2 Application
The experiments utilized the benchmark suite of vLLM v0.5.4 (rev: 4db5176) [KLZ+23].
@@ 4 Results
Conclusion 1: The average overhead is less than 7%. We quantified the overhead by measuring the throughput with TEE mode enabled versus disabled, across varying input sizes and model configurations, as shown in Table 1.
@@
The throughput is measured in two ways: the average throughput of the outputted tokens per second (TPS), and that of the parallel requests the hardware can handle (QPS). TPS is measured by running the model with a batch size of 1. It shows the pure latency overhead introduced by the TEE mode and reflects the performance of real-time requests. QPS is measured by maximizing the query throughput with a dynamically optimized batch size. It reflects the minimal average overhead the TEE mode brings.
Conclusion 2: The overhead reduces toward zero as the model size grows. As shown in Table 1, the smallest model (Llama-3.1-8B) has the highest overhead. The medium-sized model (Phi-3-14B-128k) has roughly two-thirds of the overhead compared to the smaller one. The largest model (Llama-3.1-70B) has a trivial overhead close to zero.
Conclusion 3: The latency is the main factor contributing to the overhead of the TEE mode. Table 2 shows the overhead introduced to the latency measured by TTFT and ITL. TTFT has a higher overhead compared with ITL, indicating the bottleneck is likely introduced by the I/O instead of the computation happening inside the TEE. Nevertheless, the overhead becomes trivial when hosting heavy computation models like Llama-3.1-70B.
@@
Conclusion 4: The overhead reduces as the token size grows. As shown in Figure 1, the throughput overhead reduces when the sequence length grows, measured by the total input and output token count. The detailed throughput metrics across various sequence lengths can be found in Table 3.
@@
Conclusion 5: TEE can reach typical throughput
Our experiments revealed that, with medium-sized inputs, the nVIDIA H100 GPU achieves 130 TPS for Llama-3.1-8B, while the larger Phi-3-14B model reaches approximately 6 TPS. These results demonstrate the robust performance of the H100 GPU across models of varying complexity.
@@ 5 Conclusion
Our results show that as input size grows, the efficiency of TEE mode increases significantly. When computation time within the GPU dominates overall processing time, the I/O overhead introduced by TEE mode diminishes, allowing efficiency to approach nearly 99%.
Efficiency growth is more pronounced in larger models, such as Phi3-14B-128k and Llama-3.1-70B, due to their greater computational demands, which result in longer GPU processing times. Consequently, the I/O overhead becomes increasingly trivial as model size increases.
The total token size (sum of input and output token size) significantly influences the throughput overhead. Larger total token counts lead to higher efficiencies, as they enhance the ratio of computation time to I/O time.